A particle P is moving along a straight line AB with its initial velocity 3 m/sec when

started from point A and with acceleration of 1.5 m/sec 2 . Another particle Q starts from

point A with velocity of 11.4 m/sec and acceleration of 1.2 m/sec 2 , 3 seconds after P has

left the point A. (a) determine the distance from point A when Q will overtake P, (b)

Further show that once Q will overtake P, Q will never be ahead of P by more than 9.6

m.


1
Expert's answer
2021-07-26T17:11:02-0400

Let "t=" time of motion of particle Q. Then the time of motion of particle P will be "t+3."

The equation of motion


"s=s_0+v_0t+\dfrac{at^2}{2}"

Given

"s_{0P}=0\ m,v_{0P}=3 \ m/s, a_P=1.5\ m/s^2"

"s_{0Q}=0\ m,v_{0P}=11.4 \ m/s, a_P=1.2\ m/s^2"

"s_P(t)=0+3(t+3)+\dfrac{1.5(t+3)^2}{2}"

"s_Q(t)=0+11.4t+\dfrac{1.2t^2}{2}"

(a) Q will overtake P


"3(t+3)+\dfrac{1.5(t+3)^2}{2}=11.4t+\dfrac{1.2t^2}{2}"

"3t+9+0.75t^2+4.5t+6.75-11.4t-0.6t^2=0"

"0.15t^2-3.9t+15.75=0"

"0.05t^2-1.3t+5.25=0"

"t^2-26t+105=0"




"(t-5)(t-21)=0"

"t=5"


"s_Q(5)=11.4(5)+\dfrac{1.2(5)^2}{2}=72"

"s_P(5)=3(5+3)+\dfrac{1.5(5+3)^2}{2}=72"

"t=21"


"s_Q(21)=11.4(21)+\dfrac{1.2(21)^2}{2}=504"

"s_P(21)=3(21+3)+\dfrac{1.5(21+3)^2}{2}=504"

Q will overtake P at distance of 72 m from point A.



"QP=11.4t+\dfrac{1.2t^2}{2}-(3(t+3)+\dfrac{1.5(t+3)^2}{2}),5\leq t\leq21"

"f(t)=QP=-0.15t^2+3.9t-15.75"

"t_{vertex}=-\dfrac{3.9}{2(-0.15)}=13"

"f(13)=QP_{t=13}=-0.15(13)^2+3.9(13)-15.75=9.6"

The function "f(t)" has the absolute maximum with value of "9.6" for "5\leq t\leq21."


Therefore once Q will overtake P, Q will never be ahead of P by more than 9.6 m.

P will overtake Q at "t=21" sec.